# Qubit Mapping and Routing tailored to Advanced Quantum ISAs: Not as Costly as You Think

### v1

@@ 6.3. Program-ISA-Topology co-exploration
Our evaluation also systematically explores how program patterns, ISA selection, and hardware topologies impact each other. We highlight some guidelines for optimal co-design particularly according to results achieved by Canopus (Table 4, Table 5, Table 6): ① Heavy-hex topology consistently leads to higher routing overhead across all these ISAs, compared to the seemingly more sparse 1D chain topology. This is mostly because most quantum algorithms are constructed in a subroutine-unrolling approach, naturally more friendly with chain topology. The QFT kernel detailed in Section 5.1 is a thorough good example. ② The combinatorial ISA involving both $\mathrm{CX}$-family and $\mathrm{iSWAP}$-family gates is much superior to either sole one. Specifically, either ZZPhase or SQiSW could lead to no more than 11% routing overhead reduction compared to CX, while Het results in more than 25% reduction than CX. This benefit is more significant for circuits largely containing $\mathrm{CX}/\mathrm{CZ}$ as 2Q blocks such as qec9. ③ Gate mirroring is another approach to designing powerful quantum ISAs, as both ZZPhase and SQiSW achieve comparable results to Het ISA. This is due to the same optimization opportunities brought by $\mathrm{SWAP}$ mirroring. ④ For Hamiltonian simulation programs like ising, ZZPhase is essential to improve execution performance. Therein multiple 2-local Pauli rotations equivalent to $\mathrm{XX}(\theta)$ are included. As a discrete fractional $XX(\theta)$ basis gate set, ZZPhase ISA is more suitable than other gate family selections in this case.
The real-machine experiment in Section 5.1 showcases how our method can help achieve superior compilation results and thus higher program fidelities for QFT kernels using the standard CX ISA ($\mathrm{CZ}$ on ibm_torino). However, extending this real-machine validation to alternative ISAs is currently challenging due to the scarcity of quantum processors equipped with well-calibrated heterogeneous gate sets. For instance, while IBM has proposed fractional gates, i.e., the continuous $\mathrm{ZZ}(\theta)$ gate set (ibmFractionalGates), their implementation details and calibration procedures are not publicly disclosed. To our knowledge, the $\mathrm{ZZ}(\theta)$ gates have the same duration as $\mathrm{CZ}$ on IBM’s Heron QPUs regardless of the rotation angle, and their error rates are consistently 1x-3x that of $\mathrm{CZ}$. This performance is far from the ideal assumptions of ZZPhase. Fortunately, a path forward is emerging with the recently proposed AshN gate scheme (chen2024one) that enables directly implementing any basis gates with the optimal gate durations. It is also experimentally evinced on transmon qubits by chen2025efficient, where multiple basis gates are calibrated in high fidelities that aligns with our cost model as well. This development may enable comprehensive, real-machine co-exploration of programs, ISAs, and hardware topologies in the near future.
@@ 6.4. Diverse-ISA compilation paradigms
Prior to this work, there are two major compilation paradigms targeting diverse ISAs: (1) Use the conventional compiler that operates entirely on the $\mathrm{CX}$-based circuit representation before ISA rebase. The final-stage rebase pass can usually be completed via optimal synthesis in efficient analytical or numerical computation (tucci2005introduction; peterson2022optimal; huang2023quantum; mckinney2025two). (2) Use brute-force approximate synthesis to perform structural search and numerical optimization to determine the synthesized circuit with minimal gate count (davis2019heuristics; kukliansky2023qfactor; patel2022quest). Sabre/TOQM and BQSKit are representative of these two paradigms, respectively. In our evaluation, BQSKit even underperforms the industrial-standard Sabre in most cases. Exceptionally, in terms of the circuit depth, BQSKit leads to better results than other baselines on sparse topologies (chain, heavy-hex), as its A*-based search for 2Q gate arrangement could exhibit advantages over long-range qubit routing, but this advantage does not hold for more connected topologies. Besides, the second numerical optimization based paradigm is of exponential computational complexity. For benchmarking the 216 medium-size cases in Table 5, the Rust-backend BQSKit requires on average 18 minutes to process each circuit with an Apple M3 Max CPU; in contrast, the Python-implemented Sabre requires only 17 seconds. Consequently, this second paradigm is ill-suited for compiling real-world programs, proving both ineffective and inefficient when targeting diverse ISAs (at least for discrete gate sets). Instead, although there is a gap between the conventional routing model and backend ISA properties, by means of routing-synthesis co-optimization, the first paradigm could be enhanced to bridge the gap between the routing model and backend ISA properties and thus provides a more viable path.
@@ 6.5. Runtime analysis
In our field tests for the set of benchmarks above, Canopus consistently exhibits 1x-2x runtime latency of Sabre. This result aligns with the complexity analysis in Section 4.5. Herein we specifically demonstrate the end-to-end runtime analysis for larger-scale quantum circuits. We use random quantum volume (cross2019validating) circuits generated by Qiskit for scalability benchmarking, which represent the most complicated circuit structures with dense 2Q gate arrangement. Each canonical gate within the quantum volume circuit contains unique canonical parameters as each 2Q unitary is randomly generated, thus there is no cached synthesis cost calculation for performance improvement in one pass. We select quantum volume circuits with two different widths (number of qubits), 15 and 20. We vary the depth of these circuits (qv_15 and qv_20) from 50 to 200. The largest size of these qv circuits is up to thousands of 2Q gates. As Figure 11 shows, Canopus leads to an average 1.31x (±1%) latency compared to Sabre for each benchmarked circuit. Both compilers’ latency scales linearly with circuit depth and width. If we compare the curve slopes, Canopus leads to 1.32x (1.30x) latency scaling slope compared to that of Sabre in terms of depth for qv_15 (qv_20) circuits.
@@ 7. Related Work
Qubit mapping/routing is one of the most well-explored topics of quantum compiler research, as it shares similar methodologies with instruction scheduling (codina2001unified; hennessy1983postpass) and register allocation (chaitin1982register; poletto1999linear) in classical computing.
@@
With the recent development of advanced quantum ISAs such as superconducting fractional gates (ibmFractionalGates), ion-trapped partial entangling gates (ionqPartialGates; yale2025realization), and the AshN gates (chen2024one; chen2025efficient), some works have begun exploring how to efficiently utilize these ISAs to make compiler optimizations closer to hardware characteristics. mckinney2024mirage investigates the practical performance of SQiSW ISA proposed by huang2023quantum and the synthesis capability when incorporating the basis gates’ mirrors into the ISA. Their modified Sabre algorithm offers a preliminary attempt at the collaborative gate decomposition and qubit routing approach, while the optimization opportunities considered therein are limited and the algorithmic techniques are not sophisticated. BQSKit (bqskit) and the series of works behind it (davis2019heuristics; wu2020qgo; kukliansky2023qfactor; younis2021qfast) provide a toolkit to rebase arbitrary 2Q unitaries to specific ISAs through approximate synthesis (structural search and numerical optimization) which is not computationally efficient. Approximate synthesis by BQSKit does not ensure optimal schemes for two-qubit and multi-qubit circuit synthesis. In addition, due to the lack of native compilation strategies and a rational synthesis cost model, kalloor2024quantum claims that alternative ISAs are hardly comparable to CX when evaluating quantum hardware roofline by BQSKit. As for applicability of expanded ISAs to QEC, Google’s latest theoretical (mcewen2023relaxing) and experimental (eickbusch2024demonstrating) works demonstrate the CX-iSWAP combination ISA could help suppress the fault-tolerant threshold. zhou2024halma proposes a routing-based method enhanced by CX-iSWAP for overcoming ancilla defects among surface code blocks while preserving encoded logical information, but it relies on manual design and experience.
@@ A.2. Quantum ISA and the synthesis capability
We assume the cost of one $\mathrm{CX}$ gate is $1.0$, polytopes in different colors denotes the minimal circuit cost (duration) for the coverage set if synthesized by $\mathrm{CX}$ and arbitrary 1Q gates. That is, on average, the number of $\mathrm{CX}$ gates required to synthesize arbitrary 2Q gates is $3$. In contrast, the number for SQiSW ISA is $2.21$ (huang2023quantum).
Monodromy polytope theory (peterson2020fixed) provides a framework for determining the synthesis coverage set and circuit cost (in 2Q depth) for any set of basis gates with specified costs, while the specific gate decomposition process is left to the synthesizer to complete. For the selected ISAs in Table 2 with the basis gate costs assumed in Equation 6, Figures 13, 13, 14, 15, 16 and 17 describes their coverage sets, respectively. With the enrichment of quantum ISA (e.g., combining gate families, involving mirror gates) and heterogeneous basis gate cost settings, the coverage set reveals a richer variety of convex polyhedra. That implies more optimization effects for the ISA-ware routing mechanism in Canopus.

### v3

@@ 6.3. Program-ISA-Topology co-exploration
(a) Routing overhead in terms of $C_{\mathrm{count}}$.
(b) Routing overhead in terms of $C_{\mathrm{depth}}$.
Figure 11. Routing overhead in terms of (a) $C_{\mathrm{count}}$ and (b) $C_{\mathrm{depth}}$ for different compilers across various device topologies and quantum ISAs.
Our evaluation also systematically explores how program patterns, ISA selection, and hardware topologies impact each other. We highlight some guidelines for optimal co-design particularly according to results achieved by Canopus (Table 4, Figure 11): ① Heavy-hex topology consistently leads to higher routing overhead across all these ISAs, compared to the seemingly more sparse 1D chain topology. This is mostly because most quantum algorithms are constructed in a subroutine-unrolling approach, naturally more friendly to chain topology. The QFT kernel detailed in Section 5.1 is a thorough good example. ② The combinatorial ISA involving both $\mathrm{CX}$-family and $\mathrm{iSWAP}$-family gates is much superior to either sole one. Specifically, either ZZPhase or SQiSW could lead to no more than 11% routing overhead reduction compared to CX, while Het results in more than 25% reduction than CX. This benefit is more significant for circuits largely containing $\mathrm{CX}/\mathrm{CZ}$ as 2Q blocks, such as qec9. ③ Gate mirroring is another approach to designing powerful quantum ISAs, as both ZZPhase and SQiSW achieve comparable results to Het ISA. This is due to the same optimization opportunities brought by $\mathrm{SWAP}$ mirroring. ④ For Hamiltonian simulation programs like ising, ZZPhase is essential to improve execution performance. Therein multiple 2-local Pauli rotations equivalent to $\mathrm{XX}(\theta)$ are included. As a discrete fractional $\mathrm{XX}(\theta)$ basis gate set, ZZPhase ISA is more suitable than other gate family selections in this case.
The real-machine experiment in Section 5.1 showcases how our method can help achieve superior compilation results and thus higher program fidelities for QFT kernels using the CX and ZZPhase ISAs via IBM Quantum Cloud. However, there are current practical hurdles to extending this real-machine validation to alternative ISAs—ones that arise primarily from the continued scarcity of quantum processors with well-calibrated heterogeneous gate sets. Fortunately, a path forward is emerging with the recently proposed AshN gate scheme (chen2024one) and its extended generalization (yang2025reqisc) that enable directly implementing any basis gates with the optimal gate durations. It is also experimentally demonstrated on transmon qubits by Chen et al. (chen2025efficient), where multiple basis gates are calibrated with high fidelity, which aligns with our cost model as well. This development may enable comprehensive, real-machine co-exploration of programs, ISAs, and hardware topologies in the near future.
@@ 6.4. Diverse-ISA compilation paradigms
Prior to this work, there are two major compilation paradigms targeting diverse ISAs: (1) Use the conventional compiler that operates entirely on the $\mathrm{CX}$-based circuit representation before ISA rebase. The final-stage rebase pass can usually be completed via optimal synthesis in efficient analytical or numerical computation (tucci2005introduction; peterson2022optimal; huang2023quantum; mckinney2025two). (2) Use brute-force approximate synthesis to perform structural search and numerical optimization to determine the synthesized circuit with minimal gate count (davis2019heuristics; kukliansky2023qfactor; patel2022quest). Sabre/TOQM and BQSKit are representative of these two paradigms, respectively. In our evaluation, BQSKit even underperforms the industrial-standard Sabre in most cases. Exceptionally, in terms of the circuit depth, BQSKit leads to better results than other baselines on sparse topologies (chain, heavy-hex), as its A*-based search for 2Q gate arrangement could exhibit advantages over long-range qubit routing, but this advantage does not hold for more connected topologies. Besides, the second numerical optimization based paradigm is of exponential computational complexity. For benchmarking the 216 medium-size cases, the Rust-backend BQSKit requires on average 18 minutes to process each circuit with an Apple M3 Max CPU; in contrast, the Python-implemented Sabre requires only 17 seconds. Consequently, this second paradigm is ill-suited for compiling real-world programs, proving both ineffective and inefficient when targeting diverse ISAs (at least for discrete gate sets). Instead, although there is a gap between the conventional routing model and backend ISA properties, by means of routing-synthesis co-optimization of Canopus, the first paradigm is enhanced to bridge the gap between the routing model and backend ISA properties and thus provides a more viable path.
@@ 6.5. Runtime analysis
In our field tests for the 216 cases above, Canopus consistently exhibits 1x-2x runtime latency of Sabre. This result aligns with the complexity analysis in Section 4.5. Herein we specifically demonstrate the end-to-end runtime analysis for larger-scale quantum circuits. We use random quantum volume (QV) (cross2019validating) circuits generated by Qiskit for scalability benchmarking, which represent the most complicated circuit structures with dense 2Q gate arrangement. Each canonical gate within the QV circuit contains unique canonical parameters as each 2Q unitary is randomly generated, thus there is no cached synthesis cost calculation for performance improvement in one pass. We select QV circuits with two different widths (number of qubits), 15 and 20. We vary the depth of these circuits (qv_15 and qv_20) from 50 to 200. The largest size of these qv circuits is up to thousands of 2Q gates. As Figure 12 shows, Canopus leads to an average 1.31x (±1%) latency compared to Sabre for each benchmarked circuit. Both compilers’ latency scales linearly with circuit depth and width. If we compare the curve slopes, Canopus leads to 1.32x (1.30x) latency scaling slope compared to that of Sabre in terms of depth for qv_15 (qv_20) circuits.
@@ 7. Related Work
Qubit mapping/routing is one of the most well-explored topics of quantum compiler research (zhu2025quantum), as it shares similar methodologies with instruction scheduling (codina2001unified; hennessy1983postpass) and register allocation (chaitin1982register; poletto1999linear) in classical computing.
@@
With the recent development of advanced quantum ISAs such as superconducting fractional gates (ibmFractionalGates), ion-trapped partial entangling gates (ionqPartialGates; yale2025realization), and the AshN gates (chen2024one; chen2025efficient; yang2025reqisc), some works have begun exploring how to efficiently utilize these ISAs to make compiler optimizations closer to hardware characteristics. mckinney2024mirage investigates the practical performance of SQiSW ISA proposed by huang2023quantum and the synthesis capability when incorporating the basis gates’ mirrors into the ISA. Their modified Sabre algorithm offers a preliminary attempt at the collaborative gate decomposition and qubit routing approach, while the optimization opportunities considered therein are limited and the algorithmic techniques are not sophisticated. BQSKit (bqskit) and the series of works behind it (davis2019heuristics; wu2020qgo; kukliansky2023qfactor; younis2021qfast) provide a toolkit to rebase arbitrary 2Q unitaries to specific ISAs through approximate synthesis (structural search and numerical optimization) which is not computationally efficient. Approximate synthesis by BQSKit does not ensure optimal schemes for two-qubit and multi-qubit circuit synthesis. In addition, due to the lack of native compilation strategies and a rational synthesis cost model, kalloor2024quantum claims that alternative ISAs are hardly comparable to CX when evaluating quantum hardware roofline by BQSKit. As for the applicability of expanded ISAs to QEC, Google’s latest theoretical (mcewen2023relaxing) and experimental (eickbusch2024demonstrating) works demonstrate that the CX-iSWAP combination ISA could help suppress the fault-tolerant threshold. zhou2024halma proposes a routing-based method enhanced by CX-iSWAP for overcoming ancilla defects among surface code blocks while preserving encoded logical information, but it relies on manual design and experience.
@@ A.2. Quantum ISA and the synthesis capability
We assume the cost of one $\mathrm{CX}$ gate is $1.0$. Polytopes in different colors denotes the minimal circuit cost (duration) for the coverage set if synthesized by $\mathrm{CX}$ and arbitrary 1Q gates. That is, on average, the number of $\mathrm{CX}$ gates required to synthesize arbitrary 2Q gates is $3$. In contrast, the number for SQiSW ISA is $2.21$ (huang2023quantum).
Monodromy polytope theory (peterson2020fixed) provides a framework for determining the synthesis coverage set and circuit cost (in 2Q depth) for any set of basis gates with specified costs, while the specific gate decomposition process is left to the synthesizer to complete. For the selected ISAs in Table 2 with the basis gate costs assumed in Equation 6, Figures 14, 14, 15, 16, 17 and 18 describes their coverage sets, respectively. With the enrichment of quantum ISA (e.g., combining gate families, involving mirror gates) and heterogeneous basis gate cost settings, the coverage set reveals a richer variety of convex polyhedra. That implies more optimization effects for the ISA-ware routing mechanism in Canopus.